OpenStudy (anonymous):

Use the ratio test to find the radius of convergence and interval of convergence of the power series.

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty} n^3 (x+3)^n\]

OpenStudy (anonymous):

R= the interval is [?] < x < -2

OpenStudy (abb0t):

I love the ratio test. It's the best test to use. Usually works for a lot of series! Anyways, the ratio test is: \[L = \lim_{n \rightarrow \infty }|\frac{ a_{n+1} }{ a_n }|\] if L < 1 it's absolutely convergent (hence convergent) L > 1 divergent L = 1 test fails. find another test.

OpenStudy (anonymous):

how do you find the radius ?

OpenStudy (abb0t):

Find L.

OpenStudy (abb0t):

\[|\frac{ (n+1)^3(x+3)^{n+1} }{ n^3(x+3)^n }|\]

OpenStudy (abb0t):

NOTE: \[(x+3)^n (x+3)^1\]

OpenStudy (abb0t):

Cancel out some terms. and find the limit.

OpenStudy (anonymous):

do you find the limit of n^3(x+3)^n

OpenStudy (abb0t):

No. The limit of the function you get AFTER you cross cancel some things in the fraction i gave up there.

OpenStudy (anonymous):

\[\frac{ (n+1)^3(x+3) }{ n^3 }\]

OpenStudy (anonymous):

right?

OpenStudy (abb0t):

yes.

OpenStudy (abb0t):

not take the limit as n approaches infinity.

OpenStudy (abb0t):

now*

OpenStudy (anonymous):

so the limit is 3+x

OpenStudy (anonymous):

how do i find the radius after the limit

OpenStudy (anonymous):

It means |x+3|<1 , and from this equ, you will find a range of x value that satisfy this equation. That is gonna be ur Radius of Convergence.